# Unsupervised full-field Bayesian inference of orthotropic hyperelasticity from a single biaxial test: a myocardial case study

### v1

@@ 2.2 Stochastic variational inference
In this study, we assume an a priori known constitutive law (see Eq. 27) and aim to stochastically infer the underlying parameter set $\bm{\theta}$. Leveraging Bayes’ theorem, we express both our prior beliefs about the parameter set and the influence of noisy measurements within a unified probabilistic framework.
We recast the deterministic optimization problem in (12) into Bayesian inference by treating the force residuals as the result of additive measurement noise (both for displacements and reaction forces). Let $\bm{\rho}(\bm{\theta})$ denote the global residual obtained by concatenating contributions from all load steps $t=1,\ldots,n_{t}$; for brevity, we omit explicit superscripts $(\cdot)^{t}$ when the meaning is clear from context. The residual stacks two components: (a) the internal-force equilibrium at unconstrained degrees of freedom and (b) the reaction-force consistency at fixed, i.e. instrumented, boundary segments:
$\bm{\rho}(\bm{\hat{\theta}})=\begin{bmatrix}\{f_{i}^{a}\}_{(a,i)\in\mathcal{D}^{\text{free}}}\\ \{R^{\beta}-\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a}\}_{\beta=1}^{n_{\beta}}\end{bmatrix}$ (13)
Note that $\bm{\rho}(\bm{\theta})$ implicitly depends on observed displacements through the predicted forces in $f_{i}^{a}$, and thus the material model parameters $\bm{\theta}$. We assume that these residuals are distributed according to a multivariate normal distribution with zero mean and isotropic variance, reflecting the assumption of additive Gaussian noise:
$\bm{\rho}(\bm{\hat{\theta}})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ (14)
In our case, both force and displacement measurements are affected by sensor noise. However, because displacements propagate in nonlinear fashion to the force residuals, the uncertainty introduced by the displacement noise is assumed to dominate. Under this assumption, the likelihood of the observed displacement $\hat{\bm{u}}$ and reaction force $\hat{R}^{\beta}$ data, given the parameters $\bm{\theta}$ and noise variance $\sigma^{2}$, becomes:
$\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\propto\,\exp\left(-\frac{1}{2\sigma^{2}}\|\,\bm{\rho}(\bm{\theta})\,\|^{2}\right)$ (15)
Again, for brevity, the superscript $(\cdot)^{t}$ is omitted in the subsequent discussion, although the likelihood distribution is computed as the product of the likelihood distribution for the residual computed at each independent quasi-static loading time snapshot.
Next, we place priors on the the noise variance $\sigma^{2}$ and the parameter vector $\bm{\theta}$ respectively to complete our Bayesian treatment of the inference problem. Because the variance $\sigma^{2}$ of the measurement noise is unknown and must be inferred along with the parameters, we place an inverse-gamma prior on it:
@@
Concomitantly, we enforce physical constraints on the parameters by requiring them to be positive and therefore enforce a truncated normal distribution for the prior. Here, we introduce a scaling hyperprior $V_{s}$ which allows us to decouple the overall measurement residual noise variance $\sigma$ from the underlying parameter variance.
@@
For this scaling hyperprior, we again enforce parameter positivity by setting an inverse gamma prior distribution:
@@
Combining the prior distributions (16) – (18) with our likelihood function (15) yields the posterior:
$\pi\left(\bm{\theta},\sigma^{2},v_{s}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)\,\propto\,\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\pi\left(\bm{\theta}\mid\sigma^{2},v_{s}\right)\,\pi\left(\sigma^{2}\right)\,\pi\left(v_{s}\right),$ (19)
This posterior serves as the foundation for either sampling-based inference or variational inference approaches, chosen according to model complexity and data volume. In simpler inference problems, sampling-based methods such as MCMC can often be relied on to approximate the posterior distribution. For example, Gibbs sampling has been successfully applied to two-dimensional full-field problems, where the material response could be inferred by solving a linear system of equations on a library of constitutive features (Joshi et al., 2022). In those cases, the constitutive models were potentially nonlinear in input but remained linear in their parameters, which keeps the posterior geometry relatively tractable for MCMC. In our setting, however, the constitutive models are nonlinear in both inputs and parameters, which induces strong parameter interactions and complex (often multimodal) posteriors. Together with the cost of evaluating data-heavy 3D full-field likelihoods, this makes traditional MCMC prohibitively expensive. In turn, we use variational inference to approximate the posterior distribution $\pi$ with a more tractable family of parametric guide distributions $q$. We therefore adopt stochastic variational inference (SVI) to approximate the posterior efficiently, enabling gradient-based optimization such as stochastic gradient descent and ADAM (Kingma and Welling, 2013; Kingma and Ba, 2017). SVI has been shown to substantially accelerate inference relative to MCMC (Rodrigues, 2022; Thorat et al., 2025). Additionally, we employ a mean-field variational family $q$ (Parisi, 1988; Consonni and Marin, 2007), which assumes statistical independence between the quantities of interest.
@@
where each factor is defined as:
$q\left(\bm{\theta}\right)\quad$ $=$ $\quad\mathcal{N}_{+}(\bm{\mu}^{q},\bm{\Sigma}^{q}\circ\bm{I})$ (21)
$q\left(\sigma^{2}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{\sigma}^{q},\beta_{\sigma}^{q})$ (22)
$q\left(v_{s}\right)\quad$ $=$ $\quad\mathcal{IG}(\alpha_{vs}^{q},\beta_{vs}^{q})$ (23)
and $\circ$ denotes the element-wise (Hadamard) product. The full set of variational parameters is denoted by $\bm{\Phi}$ and our hyperparameters are reported in Table 1. The hierarchical structure (Joshi et al., 2022; Thorat et al., 2025; Peirlinck et al., 2019) of our prior distributions is showcased in Figure 2.
Figure 2: Schematic of the hierarchical inference model. This hierarchical structure of the Bayesian inference model allows for simulteneous placement of prior distributions on the noise in the likelihood and the constitutive model parameters. The red and circular distributions represent hyperpriors for $V_{s}$ and $\sigma^{2}$, the lightblue circular distribution represent the prior on the material model parameters and the black and rectangular distribution represents the likelihood of the linear momentum balance. The black arrows denote the hierarchical structure dependencies. The hyperpriors are distributed according to inverse-gamma distributions. The model parameters are distributed according to a truncated multivariate normal distribution. Truncation occurs at numerical zero to enforce polyconvexity in the material model. The likelihood is assumed to be normally distributed.
To find the best approximation within this family, we minimize the Kullback-Leibler (KL) divergence (Kullback, 1968; Cover T., 1968) between the variational distribution $q\left(\bm{\Phi}\right)$ and the true posterior $\pi\left(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta}\right)$, where $\bm{\Phi}=\{\bm{\theta},v_{s},\sigma^{2}\}$:
$\mathrm{KL}\left[\pi(\bm{\Phi}\mid\hat{\bm{u}},\hat{R}^{\beta})\mid\mid q(\bm{\Phi})\right]=\mathbb{E}_{q(\bm{\Phi})}\left[-\log q\left(\bm{\Phi}\right)-\log\pi\left(\bm{\Phi}\right)-\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\mid\bm{\Phi}\right)\right]+\log\pi\left(\hat{\bm{u}},\hat{R}^{\beta}\right)$ (24)
The entire inference pipeline is implemented using NumPyro (Phan et al., 2019), which supports automatic differentiation and stochastic optimization with ADAM using the reparameterization trick (Kingma and Ba, 2017; Kingma and Welling, 2013), which alleviates the need to explicitly define the minimization procedure as was done in (Thorat et al., 2025). We initialize the optimization method with a learning rate of $r_{0}=0.01$ and use exponential learning rate decay $\gamma$. The learning rate is defined for each epoch $i$ as $r_{i}=r_{0}\gamma^{i/n_{T}}$.
@@
  Initial scale parameter for $v_{s}$ $\alpha_{vs}$ $10^{4}$
  Initial rate parameter for $v_{s}$ $\beta_{vs}$ $10^{4}$
@@
  Exponential scaling parameters $_{\star} [][-] b,b$$b_{f},b_{n},b_{fs}$ $\{7.474,22.063,34.802,5.691\}$
@@ 3.1 Myofiber architecture and tissue slab slicing
Informed by prior histological characterizations of myocardial tissue, we model our studied specimens as a layered orthotropic material in which cardiomyocytes align along main myofiber directions, packed by endomysial collagen into sheets, and connected via perimysial collagen (Rohmer et al., 2007; Holzapfel and Ogden, 2009; Alexander J. Wilson, 2022). This organization defines three locally orthogonal directions: the fiber direction $\bm{f}$, the sheet direction $\bm{s}$, and the sheet-normal direction $\bm{n}=\bm{f}\times\bm{s}$. To investigate the role of heterogeneity in unsupervised constitutive parameter identifiability, we simulate myocardial tissue slabs that exhibit both microstructure and geometry induced heterogeneity.
@@
Leveraging the intrinsic spatially varying myofiber architecture in the human heart (Lombaert et al., 2012), we extract three microstructurally distinct synthetic tissue slabs from the left ventricular free wall. All samples are assumed to be $10\times 10\times 1$ mm3 in size, corresponding to representative slabs from a healthy left ventricular wall which is at least 10 mm thick. The sample geometry and applied boundary conditions are also visualized in Figure 1 - left. Figure 3 showcases the microstructural variability across the slab domain, with directional fiber $\bm{f}$, sheet $\bm{s}$ and normal $\bm{n}$ vectors discretized element-wise, depend on how the tissue slicing and extraction procedure is performed:
@@
    circ-long: sections aligned with the ventricular circumferential–longitudinal plane. The in-plane $\bm{f}$ orientation and through-thickness $\bm{s}$ direction enable near-planar fiber alignment. We consider both aligned and off-axis variants depending on their angular deviation from the square edge. This slab specimen type aligns with the biaxial tension testing slabs studied by Sommer et al. (2015).
@@
    rad-heli: samples based on a radial-helical cut. Given the natural transmural variation of the myofiber angle $\bm{f}$ from $60^{\circ}$ in the sub-epicardium to $60^{\circ}$ in the sub-endocardium (Lombaert et al., 2012), this configuration places all three microstructural directions ($\bm{f},\bm{s},\bm{n}$) approximately within the sample plane by rotating the sample an additional $45^{\circ}$ along the radial axis. The resultant fiber orientation ranges from $15^{\circ}$ up to $105^{\circ}$, aligning the $\bm{f}$ fiber in the sample plane near the bottom edge of the sample and aligning the $\bm{n}$ fiber in the sample plane near the top edge.
To investigate the influence of geometric heterogeneity, we introduce a second set of specimens with a central circular occlusion with a radius of $1$ mm, see Figure 1 - right. These heterogeneous samples introduce additional strain concentration zones that we hypothesize may improve the identifiability of model parameters under limited data conditions, e.g. a biaxial tensile test to characterize orthotropic material behavior.
@@ 3.2 Biaxial loading emulation
We simulate a digital volume correlation experiment by modeling a biaxial loading protocol on the microstructurally and geometrically homogeneous and heterogeneous tissue slabs described above using the finite element method (FEM) implemented in ABAQUS/Standard (Dassault Systèmes Simulia Corp., 2025), see for example (Peirlinck et al., 2018). Our virtual tissue slabs are subjected to displacement-controlled symmetric biaxial loading, with loading ratios $\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$ based on stretch. The maximum stretch applied to the samples is $15\%$ their initial length, which relates to maximal values used in experiments (Sommer et al., 2015). We use linear tetrahedral elements to discretize the domain and record nodal displacements along with reaction forces for a sequence of $n_{t}$ load steps (see Figure 1)
Leveraging the local volume change of deformation $J=\det\left(\bm{F}\right)$, we multiplicatively decompose the deformation gradient $\bm{F}$ (5) into its volumetric $\bm{F}^{\rm{vol}}$ and isochoric $\bar{\bm{F}}$ parts (Flory, 1961),
@@
and introduce the following deformation (pseudo-)invariants (Spencer, 1984; Menzel, 2004):
@@
where $\bm{f}^{0}$, $\bm{s}^{0}$, $\bm{n}^{0}$ denote the local microstructure fiber, sheet, and normal unit vector orientations in the undeformed reference configuration. Sometimes these invariants are centered around zero, in which case the following notation is used: $K_{1}=I_{1}-3,K_{4ff}=I_{4ff}-1,K_{4nn}=I_{4nn}-1$. We adopt the orthotropic, compressible Holzapfel-Ogden (HO) material model (Holzapfel and Ogden, 2009; Guan et al., 2018) using the following strain energy density function:
@@
to compute the First Piola Kirchoff stress tensor $\bm{P}$ following (7). This variation of the HO-model was implemented using the universal material subroutine (Peirlinck et al., 2024). An analytical form of the stress tensor $\bm{P}$ is provided in B.
@@
In (27), $\bm{\theta}=\{a,b,a_{f},b_{f},a_{n},b_{n},a_{fs},b_{fs},K\}$ (see Table 1) denotes the to-be-inferred ground-truth material parameters. The Macaulay brackets $\langle x\rangle$ ensure that fiber and sheet term strain energy contributions activate only under tension (Wittrick, 1965; Wriggers et al., 2016; Peirlinck et al., 2024). Lastly, a bulk modulus is chosen based on the distribution of $J$ in the circ-long-hom samples and common compressibility considerations of the myocardium (McEvoy et al., 2018; Yin et al., 1996).
@@
Figure 4: Deformation profiles of geometrically homogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches are shown for circ-long-hom (top row), circ-rad-hom (middle row), and rad-heli-hom (bottom row) specimens subjected to biaxial loading protocols ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$) up to $\lambda^{\text{max}}=1.15$. Rightmost columns display distributions of invariants $I_{1}-3$, $I_{4ff}-1$, $I_{4nn}-1$, and $I_{4fs}$ for each specimen type respectively using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$. The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
@@
Figure 4 presents the deformation response of geometrically homogeneous slabs with varying microstructural architectures. The qualitative results of the maximum principal stretch are shown for all loading conditions and homogeneous tissue samples. On the right side of the figure we show kernel density estimations of the biaxially activated invariants - truncated at $20$ to visualize the spread of these invariants. The circ-long-hom configuration represents a classic homogenized myocardial sample, as excised by (Sommer et al., 2015), with assumed uniform fiber-sheet-normal orientations. Under biaxial loading, this results in a uniform deformation pattern with isotropic invariants $I_{1}$ ranging from $3.158$ to $3.257$ across the three loading protocols. Anisotropic contributions are moderate, with $I_{4ff}$ and $I_{4nn}$ spanning ranges from $1.025$ to $1.139$ and $1.024$ to $1.144$ respectively. Fiber-sheet shear contributions remain negligible, with $I_{4fs}$ limited to 0 up to machine precision. Introducing transmural fiber variation in the circ-rad-hom slab increases local deformation heterogeneity. The isotropic invariant $I_{1}$ expands to a broader range of $3.160$ to $3.367$, while $I_{4ff}$ and $I_{4nn}$ reach $1.194$ and $1.252$, respectively. Notably, fiber-sheet shear contributions increase, with $I_{4fs}$ reaching $0.099$. In the rad-heli-hom specimen, which introduces a more asymmetric helical fiber distribution, deformation heterogeneity is further amplified. The isotropic invariant $I_{1}$ ranges from $3.147$ to $3.359$, while $I_{4ff}$ increases significantly, spanning up to $1.248$. The $I_{4nn}$ component reaches $1.251$, and fiber-sheet shear contributions grow to $0.114$. The distributions of invariants shown in the right column of Figure 4 emphasize these trends. The circ-long-hom specimen exhibits single homogeneous singular values for $I_{4ff}$ and $I_{4nn}$, while $I_{4fs}$ remains zero. In contrast, circ-rad-hom and rad-heli-hom configurations display broader distributions across all invariants, particularly for fiber-sheet shear deformations. Consistent with these observations, we note tissue bulging effects in the thickness direction. Circ-rad configurations exhibit bulging centrally, whereas rad-heli configurations deform near the specimen edges. In particular, the rad-heli-hom configuration shows noticeable thickness bulging coinciding with bands of reduced strain, where fibers align in-plane. These bands are visually distinct due to their contrasting strain magnitudes relative to surrounding regions. These results underline the capacity of microstructural heterogeneity alone, even in geometrically simple specimens, to enrich deformation diversity. Such diversity is critical for reliably inferring nonlinear orthotropic constitutive behavior, motivating their role as training configurations in this study.
@@
Figure 5: Deformation profiles of geometrically heterogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches circ-long-het (top row), circ-rad-het (middle row), and rad-heli-het (bottom row) specimens under biaxial loading ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$). Invariant distributions of $I_{1}-3$, $\bar{I}_{4ff}-1$, $\bar{I}_{4nn}-1$, and $I_{4fs}$ are shown in the rightmost panels using approximated kernel density estimations denoted as $\hat{\pi}(I_{*})$. The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
@@
4 Orthotropic constitutive parameter inference

### v2

@@ 2.2 Stochastic variational inference
In this study, we assume an a priori known constitutive law (28) and aim to infer the underlying constitutive parameter set $\bm{\theta}$ together with its uncertainty. Leveraging Bayes’ theorem, we combine prior beliefs about $\bm{\theta}$ with a likelihood model that accounts for the impact of measurement noise on the physics-based residual constraints conditioned on the displacement and reaction force data.
We recast the deterministic optimization problem in (12) into Bayesian inference by treating the equilibrium residuals as the result of additive measurement noise. For a given snapshot $t=1,\ldots,n_{t}$, we define the residual vector
$\bm{\rho}^{t}(\bm{\theta})=\begin{bmatrix}\left\{f_{i}^{a,t}\right\}_{(a,i)\in\mathcal{D}^{\text{free}}}\\ \left\{R^{\beta,t}-\sum_{(a,i)\in\mathcal{D}^{\text{fix}}_{\beta}}f_{i}^{a,t}\right\}_{\beta=1}^{n_{\beta}}\end{bmatrix}$, (13)
and stack all snapshots into a single global residual
$\bm{\rho}(\bm{\theta})=\left[\begin{array}[]{c}\bm{\rho}^{1}(\bm{\theta})\\ \vdots\\ \bm{\rho}^{n_{t}}(\bm{\theta})\end{array}\right]$ (14)
This construction removes the need to carry superscript $(\cdot)^{t}$ in the subsequent probabilistic expressions while still retaining the multi-snapshot information. Note that $\bm{\rho}(\bm{\theta})$ implicitly depends on observed displacements and reaction forces through the constitutive response leading to the predicted forces in $f_{i}^{a}$.
We assume that the stacked residuals follow a multivariate normal distribution with zero mean and isotropic variance, reflecting the assumption of additive Gaussian noise:
$\bm{\rho}(\bm{\theta})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})$ (15)
which yields the likelihood
$\pi\left(\bm{u},R^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\propto\,\exp\left(-\frac{1}{2\sigma^{2}}\|\,\bm{\rho}(\bm{\theta})\,\|^{2}\right)$ (16)
In our case, both displacement field and reaction force measurements are affected by sensor noise, and we treat $\sigma^{2}$ as an effective residual variance that captures this combined uncertainty. Given that displacement propagates in a nonlinear fashion to the force residuals, we expect the residual variance to be dominated by displacement noise.
Next, we place priors on the residual noise variance $\sigma^{2}$ and the parameter vector $\bm{\theta}$. Since the variance of the measurement noise $\sigma^{2}$ is unknown and must be inferred, we assign an inverse-gamma prior:
@@
Concomitantly, we enforce physical constraints on the parameters by requiring them to be positive and therefore enforce a truncated normal distribution for the prior (truncated on the left at zero). Here, we introduce a scaling hyperprior $v_{s}$ to decouple the overall measurement residual noise variance $\sigma^{2}$ from the underlying parameter prior variance:
@@
We place an inverse-gamma hyperprior on $v_{s}$:
@@
The hierarchical structure [32, 53, 54] of our prior distributions is showcased in Figure 2.
Figure 2: Schematic of the hierarchical inference model. This hierarchical Bayesian model places priors on both the residual variance in the likelihood and the constitutive model parameters. The red circular distributions represent hyperpriors for $v_{s}$ and $\sigma^{2}$, the light-blue circular distribution represents the prior on the material model parameters, and the black rectangular distribution represents the likelihood induced by the linear momentum balance residuals. The black arrows denote hierarchical dependencies. The hyperpriors are inverse-gamma distributed, the consitutive parameters are distributed according to a truncated multivariate normal distribution, and the likelihood is assumed to be normally distributed.
Combining the prior distributions (17)–(19) with the likelihood (16) yields the posterior:
$\pi\left(\bm{\theta},\sigma^{2},v_{s}\mid\bm{u},R^{\beta}\right)\,\propto\,\pi\left(\bm{u},R^{\beta}\mid\bm{\theta},\sigma^{2}\right)\,\pi\left(\bm{\theta}\mid\sigma^{2},v_{s}\right)\,\pi\left(\sigma^{2}\right)\,\pi\left(v_{s}\right),$ (20)
This posterior can, in principle, be approximated using sampling-based inference or variational inference. In simpler inference problems, sampling-based methods such as MCMC can often be relied on to approximate the posterior distribution. For example, Gibbs sampling has been successfully applied to two-dimensional full-field problems, where the material response could be inferred by solving a linear system of equations on a library of constitutive features [32]. In those cases, the constitutive models were potentially nonlinear in input but remained linear in their parameters, which keeps the posterior geometry relatively tractable for MCMC. In our setting, however, the constitutive models are nonlinear in both inputs and parameters, which induces strong parameter interactions and complex (often multimodal) posteriors. Together with the cost of evaluating data-heavy 3D full-field likelihoods, this makes conventional MCMC-based posterior exploration computationally unattractive in practice.
We therefore adopt stochastic variational inference (SVI) to approximate the posterior efficiently using gradient-based optimization. SVI has been shown to substantially accelerate inference relative to MCMC in related settings [55, 53]. In SVI, we approximate the posterior distribution $\pi$ with a tractable family of parametric guide distributions $q$. We employ a mean-field variational family $q$ [56, 57], which assumes statistical independence between the quantities of interest,
@@
with factors:
$q\left(\bm{\theta}\right)\quad$ $=$ $\quad\mathcal{N}_{+}\left(\bm{\mu}^{q},\bm{\Sigma}^{q}\circ\bm{I}\right)$ (22)
$q\left(\sigma^{2}\right)\quad$ $=$ $\quad\mathcal{IG}\left(\alpha_{\sigma}^{q},\beta_{\sigma}^{q}\right)$ (23)
$q\left(v_{s}\right)\quad$ $=$ $\quad\mathcal{IG}\left(\alpha_{vs}^{q},\beta_{vs}^{q}\right),$ (24)
where $\circ$ denotes the element-wise Hadamard product. The full set of variational parameters is denoted by $\bm{\Phi}=\left\{\bm{\mu}^{q},\bm{\Sigma}^{q},\alpha_{\sigma}^{q},\beta_{\sigma}^{q},\alpha_{vs}^{q},\beta_{vs}^{q}\right\}$, and hyperparameters are reported in Table 1. We specifically highlight $\bm{\mu}^{q}$, and $\bm{\Sigma}^{q}$ to be the location and scale parameters of the truncated normal distribution. These correspond to the mean and standard deviation of the underlying and untruncated normal distribution used in the definition of $\mathcal{N}_{+}$.
To find the best approximation within this family, we minimize the Kullback-Leibler divergence [58, 59] between the variational distribution $q$ and the true posterior $\pi$, i.e.
$\bm{\Phi}^{\star}\leftarrow\arg\min_{{\bm{\Phi}}}\operatorname{KL}\left(q_{\bm{\Phi}}\left(\bm{\theta},v_{s},\sigma^{2}\right)\|\,\pi\left(\bm{\theta},v_{s},\sigma^{2}\mid\bm{u},R^{\beta}\right)\right),$ (25)
The entire inference pipeline is implemented in NumPyro [60], enabling automatic differentiation and stochastic optimization with ADAM using the reparameterization trick [61, 62]. Doing so, we alleviate the need to explicitly define the minimization procedure as was done in Thorat et al. [53]. We initialize the optimizer with a learning rate of $r_{0}=0.01$, and apply an exponential learning rate decay $\gamma$, i.e. $r_{i}=r_{0}\gamma^{i/n_{T}}$ at epoch i with $n_{T}$ denoting the total number of epochs.
@@
  Initial scale parameter for $v_{s}$ $\alpha_{v_{s}}$ $10^{4}$
  Initial rate parameter for $v_{s}$ $\beta_{v_{s}}$ $10^{4}$
@@
  Exponential scaling parameters $_{\star} [-] b,b$$b_{f},b_{n},b_{fs}$ $\{7.474,22.063,34.802,5.691\}$
@@ 3.1 Myofiber architecture and tissue slab slicing
Informed by prior histological characterizations of myocardial tissue [8, 9, 14], we model our studied specimens as a layered orthotropic material in which cardiomyocytes align along main myofiber directions, packed by endomysial collagen into sheets, and connected via perimysial collagen. This organization defines three locally orthogonal directions: the fiber direction $\bm{f}$, the sheet direction $\bm{s}$, and the sheet-normal direction $\bm{n}$. To investigate the role of heterogeneity in unsupervised constitutive parameter identifiability, we simulate myocardial tissue slabs that exhibit both microstructure-induced and geometry-induced heterogeneity.
@@
Leveraging the intrinsic spatially varying myofiber architecture in the human heart [63], we extract three microstructurally distinct synthetic tissue slabs from the left ventricular free wall. All samples are assumed to be $10\times 10\times 1$ mm3 in size, representative of slabs from a healthy human left ventricular wall. The sample geometry and applied boundary conditions are summarized in A.
Figure 3 showcases the microstructural variability across the slab domain, with directional fiber $\bm{f}$, sheet $\bm{s}$ and normal $\bm{n}$ vectors discretized element-wise, and their spatial distribution depends on the chosen slicing orientation:
@@
    circ-long: sections aligned with the ventricular circumferential–longitudinal plane. The in-plane $\bm{f}$ orientation and through-thickness $\bm{s}$ direction enable near-planar fiber alignment. We consider both aligned and off-axis variants depending on their angular deviation from the square edge. This slab specimen type aligns with the biaxial tension testing slabs studied by [11].
@@
    rad-heli: samples based on a radial-helical cut. Given the natural transmural variation of the myofiber angle $\bm{f}$ from $60^{\circ}$ in the sub-epicardium to $60^{\circ}$ in the sub-endocardium [63], this configuration places all three microstructural directions ($\bm{f},\bm{s},\bm{n}$) approximately within the sample plane by rotating the sample an additional $45^{\circ}$ along the radial axis. The resultant fiber orientation ranges from $15^{\circ}$ up to $105^{\circ}$, aligning the $\bm{f}$ fiber in the sample plane near the bottom edge of the sample and aligning the $\bm{n}$ fiber in the sample plane near the top edge.
To investigate the influence of geometric heterogeneity, we introduce a second set of specimens with a central circular occlusion with a radius of $1$ mm, see A. These samples are expected to introduce additional deformation heterogeneity that we hypothesize may improve constitutive parameter identifiability under limited data conditions, e.g., when inferring orthotropic behavior from a single biaxial tensile test.
@@ 3.2 Biaxial loading emulation
We simulate a digital volume correlation experiment by modeling a biaxial loading protocol on the microstructurally and geometrically homogeneous and heterogeneous tissue slabs described above using the finite element method (FEM) implemented in ABAQUS/Standard [64], see for example [26]. Our virtual tissue slabs are subjected to displacement-controlled symmetric biaxial loading, with loading ratios $\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$ defined in terms of the prescribed edge stretches along the two in-plane loading directions. The maximum stretch applied to the samples is $15\%$ of the initial edge length, consistent with maximal values used in experiments [11]. We use linear tetrahedral elements to discretize the domain and record nodal displacements along with reaction forces for a sequence of $n_{t}$ load steps (see Figure 1)
To define the constitutive response, we use the local volume change $J=\det\left(\bm{F}\right)$ and multiplicatively decompose the deformation gradient $\bm{F}$ (5) into its volumetric $\bm{F}^{\rm{vol}}$ and isochoric $\bar{\bm{F}}$ parts [65],
@@
We then introduce the following deformation (pseudo-)invariants [66, 67]:
@@
where $\bm{f}^{0}$, $\bm{s}^{0}$, $\bm{n}^{0}$ denote the local microstructure fiber, sheet, and normal unit vector orientations in the undeformed reference configuration. We adopt the orthotropic, compressible Holzapfel-Ogden (HO) material model [14, 15] using the following strain energy density function:
@@
to compute the First Piola Kirchhoff stress tensor $\bm{P}$ following (6). This variation of the HO-model was implemented using the recently proposed universal user material subroutine in Abaqus [52].
@@
In (28), $\bm{\theta}=\{a,b,a_{f},b_{f},a_{n},b_{n},a_{fs},b_{fs},K\}$ (see Table 1) denotes the to-be-inferred ground-truth material parameters. The Macaulay brackets $\langle x\rangle$ ensure that fiber and sheet term strain energy contributions activate only under tension [68, 69, 52]. Lastly, a bulk modulus is chosen based on the distribution of $J$ in the circ-long-hom samples and common compressibility considerations of the myocardium [70, 71] to limit volumetric strains while retaining numerical stability.
To quantify how microstructural and geometric heterogeneity enrich the deformation space relevant for inference, Figures 45 report the resulting distributions of the key invariants activated under biaxial loading, which motivates our single-test identification setup.
@@
Figure 4: Deformation profiles of geometrically homogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches are shown for circ-long-hom (top row), circ-rad-hom (middle row), and rad-heli-hom (bottom row) specimens subjected to biaxial loading protocols ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$) up to $\lambda^{\text{max}}=1.15$. Rightmost columns display distributions of invariants $I_{1}-3$, $I_{4ff}-1$, $I_{4nn}-1$, and $I_{4fs}$ for each specimen type respectively using approximated kernel density estimations denoted as $\mathrm{KDE}(I_{*})$ The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
@@
Figure 4 presents the deformation response of geometrically homogeneous slabs with varying microstructural architectures. The qualitative results of the maximum principal stretch are shown for all loading conditions and homogeneous tissue samples. On the right side of the figure we show kernel density estimations of the biaxially activated invariants, truncated at $20$ to visualize the spread of these invariants. The circ-long-hom configuration represents a classic homogenized myocardial sample, as excised by [11], with assumed uniform fiber-sheet-normal orientations. Under biaxial loading, this results in a uniform deformation pattern with isotropic invariants $I_{1}$ ranging from $3.158$ to $3.257$ across the three loading protocols. Anisotropic contributions are moderate, with $I_{4ff}$ and $I_{4nn}$ spanning ranges from $1.025$ to $1.139$ and $1.024$ to $1.144$ respectively. Fiber-sheet shear contributions remain negligible, with $I_{4fs}$ limited to $0.000$ up to machine precision. Introducing transmural fiber variation in the circ-rad-hom slab increases local deformation heterogeneity. The isotropic deformation invariant $I_{1}$ expands to a broader range of $3.160$ to $3.367$, while $I_{4ff}$ and $I_{4nn}$ reach $1.194$ and $1.252$, respectively. Notably, fiber-sheet shear contributions increase, with $I_{4fs}$ reaching up to $0.099$. In the rad-heli-hom specimen, which introduces a more asymmetric helical fiber distribution, deformation heterogeneity is further amplified. The isotropic invariant $I_{1}$ ranges from $3.147$ to $3.359$, while $I_{4ff}$ increases significantly, spanning up to $1.248$. The $I_{4nn}$ component reaches $1.251$, and fiber-sheet shear contributions grow to $0.114$. The distributions of invariants shown in the right column of Figure 4 emphasize these trends. The circ-long-hom specimen exhibits single homogeneous singular values for $I_{4ff}$ and $I_{4nn}$, while $I_{4fs}$ remains zero. In contrast, circ-rad-hom and rad-heli-hom configurations display broader distributions across all invariants, particularly for fiber-sheet shear deformations. Consistent with these observations, we note tissue bulging effects in the thickness direction. Circ-rad configurations exhibit bulging centrally, whereas rad-heli configurations deform near the specimen edges. In particular, the rad-heli-hom configuration shows noticeable thickness bulging coinciding with bands of reduced strain, where fibers align in-plane. Overall, microstructure-induced heterogeneity alone enriches the activated invariant space under a single biaxial protocol, which motivates using these slabs as training configurations in this study.
@@
Figure 5: Deformation profiles of geometrically heterogeneous tissue slabs with varying microstructural heterogeneity. Maximum principal stretches circ-long-het (top row), circ-rad-het (middle row), and rad-heli-het (bottom row) specimens under biaxial loading ($\lambda_{1}:\lambda_{2}=(1:2),(1:1),(2:1)$). Invariant distributions of $I_{1}-3$, $\bar{I}_{4ff}-1$, $\bar{I}_{4nn}-1$, and $I_{4fs}$ are shown in the rightmost panels using approximated kernel density estimations denoted as $\mathrm{KDE}(I_{*})$ .̇. denoted as $\mathrm{KDE}(I_{*})$. The kernel density estimation is truncated at an arbitrary level of 20 to show the spread of the invariants more clearly.
@@
4 Inference results and robustness to measurement noise
4.1 Noise-free inference: identifiability gains from microstructural and geometrical heterogeneity